### v6

# Pushing Photons with Electrons: Observation of the Polariton Drag Effect

Abstract
We show the direct effect of free electrons colliding with polaritons, changing their momentum. The result of this interaction of the electrons with the polaritons is a change in the angle of emission of the photons from our cavity structure. Because the experiment is a photon-in, photon-out system, this is equivalent to optical beam steering of photons using a direct electrical current. The effect is asymmetric, significantly slowing down the polaritons when they move oppositely to the electrons, while the polariton momentum only slightly increases when electrons moving in the same direction. We present a theoretical model which describes this effect.
It has long been known that an electron can impart momentum to an exciton-polariton through the scattering interaction with the exciton part of the polariton. Apart from the intrinsic interest in demonstrating this effect, it may have application in steering the direction of light emission by a direct electrical current, with the angle of deflection directly controlled by the applied current. However, observation of this effect has remained elusive. In this Letter we report observation of this effect with exciton-polaritons in a solid-state microcavity.
The physics of exciton-polaritons has been widely explored in recent years, and is well summarized in recent reviews, many of which have focused on Bose-Einstein condensation of polaritons [1, 2, 3, 4]. In the present work, we use a polariton condensate, but the effect is not fundamentally one that only occurs for a condensate; rather, the condensate produces a spectrally narrow emission over a wide area that makes the polariton drag effect easy to observe. Superfluidity of the condensate does not prevent the drag effect, because it is equivalent to a body force, or longitudinal force. The structures we use also allow long-distance transport over hundreds of microns [5, 6], which allows good momentum resolution in our measurements.
Recent theoretical work taking into account the polaron effect on an exciton interacting with a Fermi sea [7] has predicted that applied direct current will give a drag force on exciton-polarons, and that the effect will also be significant when the excitons are coupled to photons, as in an exciton-polariton system [8]. However, the drag effect can be understood even apart from the polaron effect as a purely collisional exchange of momentum due to Coulombic collisions between electrons and the excitonic part of a polariton, as discussed, e.g., in Ref. 9 and references therein. Other work [10, 11, 12] envisioned a drag effect between separated layers of excitons and free electrons, but this is not crucial for the effect, as free electrons can also move in the same quantum wells as the excitons.
Experimental Method. Exciton-polaritons in the strong-coupling limit were made by placing semiconductor quantum wells at the antinodes of a planar optical cavity. These semiconductor layers have excitons nearly resonant in energy with a cavity photon mode. If the -factor of the cavity is high enough, and the coupling of the photon and exciton states is strong enough, new eigenstates appear which are no longer purely photon or exciton, but a superposition of both. In other words, a photon in the cavity spends some fraction of its time as an excited electron-hole pair.
The sample used in this experiment consists of a microcavity formed by two distributed Bragg reflectors (DBRs), grown by molecular beam epitaxy (MBE). The DBRs were both made of alternating layers of Al0.2Ga0.8As and AlAs, with 32 periods in the top DBR and 40 in the bottom. 4 quantum wells (QWs), made of 7 nm GaAs layers with AlAs barriers, were placed at each of the 3 antinodes of the cavity. This microcavity design is the same used in previous work [5, 6, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] (see the Supplemental Material for more details of the sample characteristics). Long wires () were formed by etching away the top DBR, confining the polaritons within the wire and exposing the QWs.
To allow co-linear motion of free electrons in the same medium, two NiAuGe contacts were then placed upon the QWs at the ends of the wire, allowing electrical injection into the QWs (see Figure 1). Hall measurements show that the wire regions are intrinsically -doped of the order of cm-3, and the contacts are -type with heavy doping of the order of cm-3. A continuous-wave (CW) stabilized M Squared Ti:sapphire pump laser was used to excite a spot on the straight part of the wire, with a spot size of about 20 m FWHM, in the same arrangement used in Ref. 17. The pump was non-resonant, with an excess energy of about 100 meV, and mechanically chopped at 400 Hz with a pulse width of about s. A source meter was used to sweep the applied voltage along the wires while measuring electrical current. The details of the fabrication and the electrical measurement are discussed in Ref. 16. The cavity detuning of the wire device presented in the main paper was , corresponding to an exciton fraction for the lower polariton at . A full characterization of the sample is presented in the Supplementary Material.
Off-resonant optical pumping of this type of structure produces a cloud of excitons, which then lose energy and fall down into polariton states. As in previous experiments with similar structures [14, 17], there are two density thresholds for the optical excitation. At the lower threshold, a local quasicondensate is formed at the excitation spot, which can then ballistically expand away from the excitation region; above a higher critical threshold, the condensate jumps down dramatically into a much lower energy state, which is spectrally very narrow ( meV width at half maximum), has high coherence length ( m), and fills most of the available space in a potential-energy minimum. We define the pump power needed to reach this second threshold as . The condensate was observed by recording the photons that leak out of the top mirror, using conventional imaging optics for both real-space (near field) and momentum-space (far field, Fourier plane) images, and a spectrometer for energy resolution. The leakage of photons out of the cavity was a tiny fraction of the population at any moment in time, because the of the cavity is very high (), so that a steady-state population was maintained [23]. The long polariton lifetime of 300-400 ps allows macroscopic flow over hundreds of microns [24] and demonstration of true equilibrium, confirmed by fits of the particle occupation number as a function of energy to an equilibrium Bose-Einstein distribution [25, 26]. In this steady state, the energy and spatial distribution of the polariton condensate is determined by the external potential profile it feels, which is a combination of the single-particle polariton dispersion and the repulsion of polaritons from slow-moving excitons with much higher mass, and the density-dependent, repulsive polariton-polariton interaction, which tends to flatten any external potential felt by the polaritons.
Experimental Results. The experimental assignment of the direction of positive and relative to the electrical contacts and pump spot is shown in Figure 1(b). Figure 2 shows the polariton distribution along the wire at two different pump powers above threshold and at zero applied voltage. The energy “hill” near 75 m is due to the large population of excitons at the pump spot, which repulsively interact with the polaritons. A shallow cavity gradient along the -direction caused a small slope in the potential, with a total energy drop of about 0.5 meV across the wire. At the lower power, a large mono-energetic condensate forms on the side away from the pump, filling a region of the wire about 100 m long. At higher power, the population on the right becomes large enough to have repulsive interactions that shift it up in energy, forming a single, nearly mono-energetic condensate.
Figure 3 shows the energy vs. in-plane momentum () of the polaritons in the level region, which is on the right side of pump spot (Figure 2), at three different applied voltages. With zero applied voltage (Figure 3(b)), an overall nonzero in-plane momentum is observed, which is due to flow in the main part of the wire, away from the pump spot, in the -direction. With negative applied voltage (Figure 3(c)), while all the other experimental conditions remain the same, the overall momentum is clearly reduced. This voltage corresponds to conventional current flowing in the -direction, and thus electron current flowing in the -direction, which opposes the polariton flow. With positive applied voltage (Figure 3(a)), the overall momentum is increased. This clearly shows an effect of drag upon the polaritons from the electrical current.
Figure 3(d) shows the polariton distribution vs. for multiple applied voltages. A clear shift in the momentum is observed under applied voltages. Simple calculations indicate that the electrons move slower than the polaritons. The polariton velocity is measured directly from their momentum; 0.1 m-1 corresponds to cm/s. The electron velocity can be estimated from the mobility, which our measurements give as approximately 100 cm2/V-s. For a voltage drop of 15 V over 200 m, this gives electron velocities of the order of cm/s.
Theory. We have developed a quantum Boltzmann model showing the drag effect originating from the collisions between polartions and electrons. We consider polariton-electron scattering where the two particles scatter from the state to state through a scattering interaction . The interaction potential is given by
where and are the polariton creation and annihilation operators respectively and and are the electron creation and annihilation. The drag force that the electrons exert on a polariton with momentum is calculated using Fermi’s Golden rule
The wavevectors and correspond to the incoming and outgoing polariton with mass , respectively and and correspond to the incoming and outgoing electron with mass . The polaritons are taken to be 1-dimensional due to their light mass and the electrons are assume to be 2-dimensional. The -function accounts for the energy conservation in the polariton-electron scattering, is the occupation of the polaritons, and is the occupation of the electrons, where is the drift wavevector of the electrons induced by the voltage difference and is the drift wavevector of the polaritons induced by ballistically expanding away from the excitation region. A detailed derivation of this model is discussed in the Supplemental Material.
Figure 4(a) shows the force polaritons feel due to electron-polariton scattering as a function of the relative velocity between the electrons and polaritons, defined as . When the relative velocity is zero, the polaritons feel no drag force from the electrons. Depending on the sign of the relative velocity, the polaritons can either feel a positive force (push) or a negative force (drag) due to collisions with electrons.
To simulate the experimental data, we fit the three different steady-state curves in Fig. 3 with a drifted Bose-Einstein distribution and extracted the steady state drift momentum of the polaritons for the three different voltage cases (see Supplemental Material for more details). We then computed the average force the drifted polaritons feel for the three different voltage cases, which is defined as
We find that all three cases can be fit with the same polariton density and almost exactly the same steady state drag force magnitude, as seen in Fig. 4(b). This implies that the driving force on the polaritons, which is a combination of gradient of the exciton cloud potential and the cavity gradient, is nearly the same in all three cases. This is consistent with our experimental observation that the exciton cloud profile is not affected by the applied voltage, as seen in spatially resolved images presented in the Supplementary material. As discussed in the Supplementary material, although there are slight energy differences (ca. 20 eV) of the polariton energy in different cases, these appear as changes of the overall average energy and not changes in the gradient; these energy changes are also not correlated with the direction of the current.
Our model indicates that in the three different voltgae cases in Fig. 3, the polaritons are slowed down since the polaritons are faster than the electrons. In the case where the electrons move along the same direction as the polaritons, the drag force is less than when they move in the opposite direction, which gives the shifts of the momentum shown in Fig. 3.
Because the photoluminescence intensity, which directly indicates the density of the condensate, was not found to be significantly different for different applied voltages, we can rule out the shifts in the condensate energy due to the polariton-polariton and polariton-exciton interactions. What remains is the effect of the the electron drag in our model.
Conclusions. We have demonstrated that a direct current can directly alter the momentum of exciton-polaritons; this has the effect of changing the angle of photon emission. Since the polaritons are effectively renormalized photons, created by photon absorption and ending with photon emission, this polariton drag is, in effect, using a direct electric current to change the momentum of photons. As polariton structures move ever closer to practical room temperature devices [4], this basic effect may also be possible in those devices.
We used a simple kinetic theory to explain the effect of drag of a Bose gas interacting with a thermal, fermionic reservoir. We showed that the Bose gas experiences a drag force, which is a function of the relative velocity between the Bose gas and the fermionic reservoir. We further show in the Supplemental Material, when the superfluid velocity of the condensate exceeds the relative velocity between the Bose gas and the fermionic reservoir, the drag force decreases, but it is nonzero. This indicates that it is simplistic to say that condensates do not experience drag or dissipation. In steady state, when there is interaction with an incoherent reservoir of non-condensate particles, they experience both.
Funding: The work at Pittsburgh was funded by the Army Research Office (W911NF-15-1-0466) and by the National Science Foundation (Grant No. DMR-2004570). The work of sample fabrication at Princeton was funded in part by the Gordon and Betty Moore Foundation (GBMF-4420) and by the National Science Foundation MRSEC program through the Princeton Center for Complex Materials (DMR-0819860).
Author Contributions: The experimental work was carried out by D.M.M and Q.Y. under the supervision of D.W.S. D.M.M and Q.Y. contributed equally to this work. The theoretical model was developed by H.A. with the help of S.M. The sample was grown by L.N.P and K.W., and was designed and characterized by J.B. D.M.M and Q.Y. designed the devices. The device in the main text was fabricated by Q.Y. with the help of B.O. The device in the Supplemental Material was fabricated by B.O. D.M.M, Q.Y., H.A. and D.W.S have done most of the writing.
See Supplemental Material at [URL TBD] for characteristics of the sample and fabricated devices, additional experimental data and details about the numerical model, comparison to other works showing apparently similar effects, and additional analysis of the influence of injected electrons.

### v7

@@
 We present a theoretical model which describes this effect.
-It has long been known that an electron can impart momentum to an exciton-polariton through the scattering interaction with the exciton part of the polariton.
+It has long been known that an electron can impart momentum to an exciton-polariton through the scattering interaction with the exciton part of the polariton [1].
 Apart from the intrinsic interest in demonstrating this effect, it may have application in steering the direction of light emission by a direct electrical current, with the angle of deflection directly controlled by the applied current.
@@
 In this Letter we report observation of this effect with exciton-polaritons in a solid-state microcavity.
-The physics of exciton-polaritons has been widely explored in recent years, and is well summarized in recent reviews, many of which have focused on Bose-Einstein condensation of polaritons [1, 2, 3, 4].
+The physics of exciton-polaritons has been widely explored in recent years, and is well summarized in recent reviews, many of which have focused on Bose-Einstein condensation of polaritons [2, 3, 4, 5, 6, 7, 8, 9].
+In particular, there has been great interest in one-dimensional wires for conducting polariton transport [10, 11, 12].
 In the present work, we use a polariton condensate, but the effect is not fundamentally one that only occurs for a condensate; rather, the condensate produces a spectrally narrow emission over a wide area that makes the polariton drag effect easy to observe.
@@
 4 quantum wells (QWs), made of 7 nm GaAs layers with AlAs barriers, were placed at each of the 3 antinodes of the cavity.
-This microcavity design is the same used in previous work [5, 6, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] (see the Supplemental Material for more details of the sample characteristics).
+This microcavity design is the same used in previous work [13, 14, 21, 22, 23, 24] (see the Supplemental Material [25] for more details of the sample characteristics).
 Long wires () were formed by etching away the top DBR, confining the polaritons within the wire and exposing the QWs.
@@
 Hall measurements show that the wire regions are intrinsically -doped of the order of cm-3, and the contacts are -type with heavy doping of the order of cm-3.
-A continuous-wave (CW) stabilized M Squared Ti:sapphire pump laser was used to excite a spot on the straight part of the wire, with a spot size of about 20 m FWHM, in the same arrangement used in Ref.
-17.
+A continuous-wave (CW) stabilized M Squared Ti:sapphire pump laser was used to excite a spot on the straight part of the wire, with a spot size with diameter about 20 m, in the same arrangement used in Ref.
+23.
 The pump was non-resonant, with an excess energy of about 100 meV, and mechanically chopped at 400 Hz with a pulse width of about s.
@@
 The details of the fabrication and the electrical measurement are discussed in Ref.
-16.
+22.
 The cavity detuning of the wire device presented in the main paper was , corresponding to an exciton fraction for the lower polariton at .
-A full characterization of the sample is presented in the Supplementary Material.
+A full characterization of the sample is presented in the Supplementary Material [25].
 Off-resonant optical pumping of this type of structure produces a cloud of excitons, which then lose energy and fall down into polariton states.
@@
 The energy “hill” near 75 m is due to the large population of excitons at the pump spot, which repulsively interact with the polaritons.
-A shallow cavity gradient along the -direction caused a small slope in the potential, with a total energy drop of about 0.5 meV across the wire.
+A shallow cavity gradient along the -direction caused a small slope in the potential, with a total energy drop of about 0.5 meV across the whole wire.
 At the lower power, a large mono-energetic condensate forms on the side away from the pump, filling a region of the wire about 100 m long.
@@
 Simple calculations indicate that the electrons move slower than the polaritons.
-The polariton velocity is measured directly from their momentum; 0.1 m-1 corresponds to cm/s.
-The electron velocity can be estimated from the mobility, which our measurements give as approximately 100 cm2/V-s.
-For a voltage drop of 15 V over 200 m, this gives electron velocities of the order of cm/s.
+The polariton steady-state velocity is measured directly from their momentum; 0.087 m-1 corresponds to cm/s.
+The electron velocity can be estimated from the mobility; Hall measurements on these samples at room temperature give mobility of approximately 100 cm2/V-s.
+For a voltage drop of 15 V over 200 m, this gives electron velocities of the order of cm/s, which is expected to increase at low temperature [28], as implied by the fits to the data discussed below.
 Theory.
@@
 The interaction potential is given by
-where and are the polariton creation and annihilation operators respectively and and are the electron creation and annihilation.
+where and are the polariton creation and annihilation operators respectively and and are the electron creation and annihilation; is determined by momentum conservation.
 The drag force that the electrons exert on a polariton with momentum is calculated using Fermi’s Golden rule
 The wavevectors and correspond to the incoming and outgoing polariton with mass , respectively and and correspond to the incoming and outgoing electron with mass .
+We assume the polariton dispersion is quadratic, given by .
 The polaritons are taken to be 1-dimensional due to their light mass and the electrons are assume to be 2-dimensional.
 The -function accounts for the energy conservation in the polariton-electron scattering, is the occupation of the polaritons, and is the occupation of the electrons, where is the drift wavevector of the electrons induced by the voltage difference and is the drift wavevector of the polaritons induced by ballistically expanding away from the excitation region.
-A detailed derivation of this model is discussed in the Supplemental Material.
+A detailed derivation of this model is discussed in the Supplemental Material [25].
 Figure 4(a) shows the force polaritons feel due to electron-polariton scattering as a function of the relative velocity between the electrons and polaritons, defined as .
@@
 To simulate the experimental data, we fit the three different steady-state curves in Fig.
-3 with a drifted Bose-Einstein distribution and extracted the steady state drift momentum of the polaritons for the three different voltage cases (see Supplemental Material for more details).
+3 with a drifted Bose-Einstein distribution , and extracted the steady state drift momentum of the polaritons for the three different voltage cases (see Supplemental Material [25] for more details).
 We then computed the average force the drifted polaritons feel for the three different voltage cases, which is defined as
-We find that all three cases can be fit with the same polariton density and almost exactly the same steady state drag force magnitude, as seen in Fig.
-4(b).
-This implies that the driving force on the polaritons, which is a combination of gradient of the exciton cloud potential and the cavity gradient, is nearly the same in all three cases.
-This is consistent with our experimental observation that the exciton cloud profile is not affected by the applied voltage, as seen in spatially resolved images presented in the Supplementary material.
+As seen in Figure 4(b), the drag force is a function of the full width at half maximum (FWHM) of the polariton distribution .
+If the distribution is very broad, then it will feel nearly equal, canceling contributions from electrons moving in opposite directions.
+However, if the distribution is narrow and offset from zero momentum, as in our experimental , plotted as the dashed red line in Fig.
+4(a), there is a significant net force of the electrons on the polaritons.
+The theoretical model describes the steady-state velocity regime, which dominates the polariton dynamics.
+We note that the experimentally extracted velocity corresponds to the measured polariton steady-state velocity.
+Initially, the polaritons begin with a higher initial velocity, which is then slowed down by drag, through interactions with electrons and other quasiparticles in the system, ultimately reaching the steady-state value we observe.
+To fit the data, we assume that the driving force on the polaritons, which is a combination of gradient of the exciton cloud potential and the cavity gradient, is the same in all three cases.
+This is consistent with our experimental observation that the exciton cloud profile is not affected by the applied voltage, as seen in spatially resolved images presented in the Supplementary material [25].
 As discussed in the Supplementary material, although there are slight energy differences (ca.
 20 eV) of the polariton energy in different cases, these appear as changes of the overall average energy and not changes in the gradient; these energy changes are also not correlated with the direction of the current.
-Our model indicates that in the three different voltgae cases in Fig.
-3, the polaritons are slowed down since the polaritons are faster than the electrons.
-In the case where the electrons move along the same direction as the polaritons, the drag force is less than when they move in the opposite direction, which gives the shifts of the momentum shown in Fig.
+Using the experimentally measured polariton velocities, we then adjust the electron velocity in the two voltage-driven cases to give the same drag force as the case to cancel out the driving force in steady state, keeping the ratio of the electron velocities the same as the ratio of the measured currents in the two cases.
+The fact that we can do this self-consistently confirms that our assumptions are valid, in particular that the electron density remains constant, independent of the voltage, and the driving force is unchanged by the voltage.
+By contrast, a related, recent experiment on polariton drag [29] assumed that the main effect was due to a change of the exciton and electron density gradients.
+From our fits, the absolute value of the electron velocity is estimated as cm/s in the case of +14 V, which is consistent with expected increase of the measured mobility at room temperature, discussed above.
+Our model indicates that in all three different voltage cases in Fig.
+3, the polaritons are slowed down by interaction with the electrons, since the polaritons are faster than the electrons.
+In the case where the electrons move along the same direction as the polaritons, this drag force is less than when they move in the opposite direction, which gives the shifts of the momentum shown in Fig.
 3.
@@
 We have demonstrated that a direct current can directly alter the momentum of exciton-polaritons; this has the effect of changing the angle of photon emission.
-Since the polaritons are effectively renormalized photons, created by photon absorption and ending with photon emission, this polariton drag is, in effect, using a direct electric current to change the momentum of photons.
+Since the polaritons are effectively renormalized photons, created by photon absorption and ending with photon emission, this polariton drag is, in effect, using a direct electric current to change the momentum of photons, in the spirit of the historical Fizeau experiment of 1851 [30].
 As polariton structures move ever closer to practical room temperature devices [9], this basic effect may also be possible in those devices.
@@
 We showed that the Bose gas experiences a drag force, which is a function of the relative velocity between the Bose gas and the fermionic reservoir.
-We further show in the Supplemental Material, when the superfluid velocity of the condensate exceeds the relative velocity between the Bose gas and the fermionic reservoir, the drag force decreases, but it is nonzero.
 This indicates that it is simplistic to say that condensates do not experience drag or dissipation.
 In steady state, when there is interaction with an incoherent reservoir of non-condensate particles, they experience both.
-Funding: The work at Pittsburgh was funded by the Army Research Office (W911NF-15-1-0466) and by the National Science Foundation (Grant No.
-DMR-2004570).
+Funding: The work at Pittsburgh was funded by the Army Research Office (W911NF-15-1-0466) and by the National Science Foundation grants DMR-2004570 and DMR-2306977.
 The work of sample fabrication at Princeton was funded in part by the Gordon and Betty Moore Foundation (GBMF-4420) and by the National Science Foundation MRSEC program through the Princeton Center for Complex Materials (DMR-0819860).
@@
 with the help of S.M.
+and D.W.S.
 The sample was grown by L.N.P and K.W., and was designed and characterized by J.B.
@@
 and D.W.S have done most of the writing.
-See Supplemental Material at [URL TBD] for characteristics of the sample and fabricated devices, additional experimental data and details about the numerical model, comparison to other works showing apparently similar effects, and additional analysis of the influence of injected electrons.
+Data Availability: The data that support the findings of this article are openly available [42].